### v1

@@ Abstract
A comprehensive guidance view for cardiac interventional surgery can be provided by the real-time fusion of the intraoperative 2D images and preoperative 3D volume based on the ultrasound frame-to-volume registration. However, cardiac ultrasound images are characterized by a low signal-to-noise ratio and small differences between adjacent frames, coupled with significant dimension variations between 2D frames and 3D volumes to be registered, resulting in real-time and accurate cardiac ultrasound frame-to-volume registration being a very challenging task. This paper introduces a lightweight end-to-end Cardiac Ultrasound frame-to-volume Registration network, termed CU-Reg. Specifically, the proposed model leverages epicardium prompt-guided anatomical clues to reinforce the interaction of 2D sparse and 3D dense features, followed by a voxel-wise local-global aggregation of enhanced features, thereby boosting the cross-dimensional matching effectiveness of low-quality ultrasound modalities. We further embed an inter-frame discriminative regularization term within the hybrid supervised learning to increase the distinction between adjacent slices in the same ultrasound volume to ensure registration stability. Experimental results on the reprocessed CAMUS dataset demonstrate that our CU-Reg surpasses existing methods in terms of registration accuracy and efficiency, meeting the guidance requirements of clinical cardiac interventional surgery. Our code is available at https://github.com/LLEIHIT/CU-Reg.
@@ 1 Introduction
The ultrasound frame-to-volume registration aims to seek a transformation that optimally aligns the resampled slice from the given volume by the transformation with the 2D input image [6, 5], as shown in Fig. 1. Existing registration methods are divided into mathematical methods and deep learning-based methods. Mathematically, the registration task is usually modeled as an optimization problem [16, 13]. Although iteration-based methods can yield reasonable accuracy, they cannot meet the real-time requirements of cardiac surgical guidance due to the slow registration speed. Currently, various deep learning-based methods are widely applied to the image registration task, such as directly learning target transformations [18, 4], keypoint descriptors [15], and image similarity metrics [8]. In the field of frame-to-volume registration, Hou et al. [10] utilized a CNN-based model to predict rigid transformation of arbitrary 2D image slices from 3D volumes, but only attained an average alignment error of 7 mm on simulated MRI brain data. Yeung et al. [19] also employed a CNN to predict the position of 2D ultrasound fetal brain scans in 3D atlas space. However, the method only takes a set of images rather than image-volume pairs as input, which results in poor generalization ability of the model among individuals. For the ultrasound frame-to-volume registration, Guo et al. [6] introduced an end-to-end registration network to align a 2D TRUS frame with a 3D TRUS volume. However, this method extracts features from ultrasound images only using 2D and 3D convolutions and directly concatenates them, which fails to sufficiently provide critical anatomical cues and adequate cross-dimensional feature interactions for the registration of ultrasound samples with low signal-to-noise ratios.
@@
In this paper, we aim to accomplish a real-time and accurate cardiac ultrasound frame-to-volume registration to provide a complete guidance view for cardiac interventional surgery under the beating heart. To address the feature extraction difficulties caused by low signal-to-noise ratios and low tissue contrast in ultrasound images, we introduce epicardium mask prompts to provide sufficient critical anatomical information. Specifically, a bi-directional prompt-guided gated cross-dimensional attention (PGCA) operation is introduced to produce abundant structure features and perform efficient interaction between 2D frame and 3D volume features. Further, we propose a voxel-wise local-global aggregation (VLGA) module to integrate dense local-global features across dimensions. To avoid the large registration errors caused by small differences between adjacent frames, we embed an inter-frame discriminative regularization term within our hybrid supervised learning to increase the distinction between adjacent slices in the same ultrasound volume to ensure registration stability. Additionally, we build a simulated cardiac frame-to-volume registration dataset through post-processing the CAMUS dataset [12]. The experimental results demonstrate that our model achieves superior performance compared to the state-of-the-art methods, e.g., a DistErr of 3.91 mm and a runtime of over 35 FPS. We hope our model can be applied for real-time and accurate guidance in cardiac interventions.
@@ 2.1 Overview of the proposed CU-Reg
Given the moving image $U_{v}\in\mathbb{R}^{D\times H\times W}$ and fixed images $U_{f}^{i}\in\mathbb{R}^{H\times W},i=\{0,1,2,3\}$, where $U_{f}^{0}$ denotes the current anchor frame to be estimated and $U_{f}^{1,2,3}$ denotes adjacent frames within the same volume. We initially utilize two independent encoding branches to extract 2D slice features $\mathcal{F}_{s}$ and 3D volume features $\mathcal{F}_{v}$. For the 2D frame branch, we first normalize the channel dimensions using a convolution layer and feed frames into the CNN-based encoder [20] for multi-level features. To encode the 3D volume, we employ three 3D convolutional blocks with different kernel sizes to extract coarse-to-fine multi-scale features. Subsequently, we introduced the epicardium prompt-guided cross-dimensional attention operation that leverages the epicardium mask prompt with the bi-directional gated cross-dimensional attention block to spotlight critical anatomical features, providing informative alignment cues for ultrasound images. The enhanced features are processed by our voxel-wise local-global aggregation module to boost the fine-grained fusion of cross-dimensional representations. Finally, transformation parameters are estimated via the pose predictor. Additionally, our model embeds an inter-frame discriminative regularization term to highlight the discrimination between adjacent slices within the same ultrasound volume, yielding a hybrid-supervised training strategy to ensure registration stability.
@@ 2.3 Voxel-wise Dense Local-Global Aggregation
After obtaining the sufficient interaction between ultrasound frames and volume features, it is essential to fuse cross-dimensional features for a cohesive synthesis of critical structural details. To accommodate ultrasound registration with multiple noises, we introduce a voxel-wise local-global aggregation module (VLGA) to efficiently associate local dense cues with global geometric information. Given the enhanced 2D slice features $z_{s}\in\mathbb{R}^{d\times DHW}$ and 3D volume features $z_{v}\in\mathbb{R}^{d\times DHW}$ derived from bi-directional PGCA operations, we associate the volume feature of each voxel with its spatial corresponding slice feature to generate voxel-wise pairs of features. Subsequently, these feature pairs are concatenated and fed into an MLP to obtain a global feature vector $\mathcal{F}_{glo}$. Lastly, $\mathcal{F}_{glo}$ is concatenated with the paired features, facilitating the acquisition of local-global context insights. Our VLGA module can be summarized as follows:
@@
2.4 Hybird Supervised Learning
In the training phase, we employ a hybrid loss function to supervise our model. Unlike existing methods that jointly regress the pose parameters, we propose to predict the pose parameters separately by decoupling the rotation and translation branches so as to avoid discontinuity in the rotational space from disturbing the prediction of translation parameters. Here, we use two MLP layers to regress the rotation and translation parameters along with a smoothed L1 loss to supervise the pose parameters. The translation loss $\mathcal{L}_{trans}$ and rotation loss $\mathcal{L}_{rot}$ can be formulated as
@@ 2.4 Hybird Supervised Learning
where $D_{f}\in\mathbb{R}^{1\times 3}$ is the estimated inter-frame distance, which is predicted from the aggregated feature $Z$ by a separated MLP network, and $D^{*}_{f}$ is the ground truth. In addition, we leverage the MS-SSIM loss $\mathcal{L}_{sim}$ for self-supervised training by constraining the similarity between the resampled and input frames. Overall, the hybrid loss function of CU-Reg is computed as
@@ 2.5 Implementation Details
Our model is trained by an Adam optimizer on a single RTX3090 GPU for 500 epochs with a batch size of 16. The inputs of CU-Reg are ultrasound slices with the size of 128$\times$128$\times$1$\times$1 and corresponding volumes with the size of 128$\times$128$\times$32$\times$1. In the 2D ultrasound frame branch, we employ a ResNet-34 [9] as the 2D encoder. In the training phase, the inputs are a current anchor frame and three adjacent frames, i.e., 128$\times$128$\times$1$\times$4. During inference, we manually adjust the input dimension by repeating the number of channels four times.
@@ 3.1 Dataset and Evaluation Metrics
To evaluate the cardiac frame-to-volume registration network, a simulated dataset is generated by post-processing the public CAMUS dataset [12]. CAMUS contains 3D echocardiographic volumes with two- and four-chamber views of 500 patients, along with the masks of the left ventricular epicardium. Each original volume is first randomly generated with four transformations. Compared with the identity transformation, the deviations of translation parameters of each transformation are within the range of 10 mm, and the deviations of the rotation parameters are within the range of 20 degrees. Based on these transformations, 2D slices and corresponding masks are sampled from the original volume. The sampled slices include 128$\times$128 pixels, and the pixel spacing is 0.62 mm×$\times$0.62 mm. Meanwhile, a new volume is sampled from the original volume based on the identity transformation with a volume size of 128$\times$128$\times$32 and a voxel spacing of 0.62 mm×$\times$0.62 mm×$\times$0.62 mm. In this way, we can obtain four volume-frame-mask pairs with true transformations for one original volume. All data is split according to individual patients, with 3,600 volume-frame-mask pairs for training and 400 volume-frame-mask pairs for testing.
@@ 4.0.1 Acknowledgements
This work was supported in part by a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project Reference Number: T45-401/22-N), in part by an Innovation and Technology Fund under Guangdong-Hong Kong Technology Cooperation Funding Scheme (project no. GHP/050/20SZ).
@@
The authors have no competing interests to declare that are relevant to the content of this article.

### v2

@@ Abstract
Real-time fusion of intraoperative 2D ultrasound images and preoperative 3D ultrasound volume based on the frame-to-volume registration can provide a comprehensive guidance view for cardiac interventional surgery. However, cardiac ultrasound images are characterized by a low signal-to-noise ratio and small differences between adjacent frames, coupled with significant dimension variations between 2D frames and 3D volumes to be registered, resulting in real-time and accurate cardiac ultrasound frame-to-volume registration being a very challenging task. This paper introduces a lightweight end-to-end Cardiac Ultrasound frame-to-volume Registration network, termed CU-Reg. Specifically, the proposed model leverages epicardium prompt-guided anatomical clues to reinforce the interaction of 2D sparse and 3D dense features, followed by a voxel-wise local-global aggregation of enhanced features, thereby boosting the cross-dimensional matching effectiveness of low-quality ultrasound modalities. We further embed an inter-frame discriminative regularization term within the hybrid supervised learning to increase the distinction between adjacent slices in the same ultrasound volume to ensure registration stability. Experimental results on the reprocessed CAMUS dataset demonstrate that our CU-Reg surpasses existing methods in terms of registration accuracy and efficiency, meeting the guidance requirements of clinical cardiac interventional surgery. Our code is available at https://github.com/LLEIHIT/CU-Reg.
@@ 1 Introduction
The ultrasound frame-to-volume registration aims to seek a transformation that optimally aligns the resampled slice from the given volume by the transformation with the 2D input image [6, 5], as shown in Fig. 1. Existing registration methods are divided into mathematical methods and deep learning-based methods. Mathematically, the registration task is usually modeled as an optimization problem [17, 14]. Although iteration-based methods can yield reasonable accuracy, they cannot meet the real-time requirements of cardiac surgical guidance due to the slow registration speed. Currently, various deep learning-based methods are widely applied to the image registration task, such as directly learning target transformations [19, 4], keypoint descriptors [16], and image similarity metrics [8]. In the field of frame-to-volume registration, Hou et al. [10] utilized a CNN-based model to predict rigid transformation of arbitrary 2D image slices from 3D volumes, but only attained an average alignment error of 7 mm on simulated MRI brain data. Yeung et al. [20] also employed a CNN to predict the position of 2D ultrasound fetal brain scans in 3D atlas space. However, the method only takes a set of images rather than image-volume pairs as input, which results in poor generalization ability of the model among individuals. For the ultrasound frame-to-volume registration, Guo et al. [6] introduced an end-to-end registration network to align a 2D TRUS frame with a 3D TRUS volume. However, this method extracts features from ultrasound images only using 2D and 3D convolutions and directly concatenates them, which can be further enhanced by epicardium mask prompts to provide sufficient critical anatomical cues and adequate cross-dimensional feature interactions for the registration of ultrasound samples with low signal-to-noise ratios.
@@
In this paper, we aim to accomplish a real-time and accurate cardiac ultrasound frame-to-volume registration to provide a complete guidance view for cardiac interventional surgery under the beating heart. To address the feature extraction difficulties caused by low signal-to-noise ratios and low tissue contrast in ultrasound images, we introduce epicardium mask prompts to provide sufficient critical anatomical information. Specifically, a bi-directional prompt-guided gated cross-dimensional attention (PGCA) operation is introduced to produce abundant structure features and perform efficient interaction between 2D frame and 3D volume features. Further, we propose a voxel-wise local-global aggregation (VLGA) module to efficiently integrate dense local-global features across dimensions. To avoid the large registration errors caused by small differences between adjacent frame images, we embed an inter-frame discriminative regularization term within our hybrid supervised learning to increase the distinction between adjacent slices in the same ultrasound volume to ensure registration stability. Additionally, we build a simulated cardiac frame-to-volume registration dataset through post-processing the CAMUS dataset [13]. The experimental results demonstrate that our model achieves superior performance compared to the state-of-the-art methods, e.g., a runtime of over 35 FPS and a DistErr of 3.91 mm, which can meet the 5 mm accuracy requirement for many cardiac catheterizations [12]. We hope our model can be applied for real-time and accurate guidance in cardiac interventions.
@@ 2.1 Overview of the proposed CU-Reg
Given the moving image $U_{v}\in\mathbb{R}^{D\times H\times W}$ and fixed images $U_{f}^{i}\in\mathbb{R}^{H\times W},i=\{0,1,2,3\}$, where $U_{f}^{0}$ denotes the current anchor frame to be estimated and $U_{f}^{1,2,3}$ denotes adjacent frames within the same volume. We initially utilize two independent encoding branches to extract 2D slice features $\mathcal{F}_{s}$ and 3D volume features $\mathcal{F}_{v}$. For the 2D frame branch, we first use two $Conv1d$ layers to normalize the channel dimension by increasing the number of channels from 4 to 64 followed by reducing it to 3, then feed frames into the CNN-based encoder [21] for multi-level features. To encode the 3D volume, we employ three 3D convolutional blocks with different kernel sizes to extract coarse-to-fine multi-scale features. Subsequently, we introduced the epicardium prompt-guided cross-dimensional attention operation that leverages the epicardium mask prompt with the bi-directional gated cross-dimensional attention block to spotlight critical anatomical features, providing informative alignment cues for ultrasound images. The enhanced features are processed by our voxel-wise local-global aggregation module to boost the fine-grained fusion of cross-dimensional representations. Finally, transformation parameters are estimated via the pose predictor. Additionally, our model embeds an inter-frame discriminative regularization term to highlight the discrimination between adjacent slices within the same ultrasound volume, yielding a hybrid-supervised training strategy to ensure registration stability.
@@ 2.3 Voxel-wise Dense Local-Global Aggregation
After obtaining the sufficient interaction between ultrasound frames and volume features, it is essential to fuse cross-dimensional features for a cohesive synthesis of critical structural details. To accommodate ultrasound registration with multiple noises, we introduce a voxel-wise local-global aggregation module (VLGA) to efficiently associate local dense cues with global geometric information. Given the enhanced 2D slice features $z_{s}\in\mathbb{R}^{d\times DHW}$ and 3D volume features $z_{v}\in\mathbb{R}^{d\times DHW}$ derived from bi-directional PGCA operations, we map the volume feature of each voxel and its spatial corresponding slice feature to the same size through 3D convolution and 2D convolution operations to generate voxel-wise pairs of features. Subsequently, these feature pairs are concatenated and fed into an MLP to obtain a global feature vector $\mathcal{F}_{glo}$. Lastly, $\mathcal{F}_{glo}$ is concatenated with the paired features, facilitating the acquisition of local-global context insights. Our VLGA module can be summarized as follows:
@@
2.4 Hybrid Supervised Learning
In the training phase, we employ a hybrid loss function to supervise our model. Unlike existing methods that jointly regress the pose parameters, we propose to predict the pose parameters separately by decoupling the rotation and translation branches so as to avoid discontinuity in the rotational space from disturbing the prediction of translation parameters. Here, we use two MLP layers to regress the rotation and translation parameters along with a smoothed L1 loss to supervise the pose parameters. The translation loss $\mathcal{L}_{trans}$ and rotation loss $\mathcal{L}_{rot}$ are used to make the network converge quickly, and they can be formulated as
@@ 2.4 Hybrid Supervised Learning
where $D_{f}\in\mathbb{R}^{1\times 3}$ is the estimated inter-frame distance, which is predicted from the aggregated feature $Z$ by a separated MLP network, and $D^{*}_{f}$ is the ground truth. The inter-frame distance is defined as the Euclidean distance between the translation vectors of the two adjacent frames. In addition, we leverage the MS-SSIM loss $\mathcal{L}_{sim}$ for self-supervised training by constraining the similarity between the resampled and input frames to make the training process more stable and avoid overfitting. Overall, the hybrid loss function of CU-Reg is computed as
@@ 2.5 Implementation Details
Our model is trained by an Adam optimizer on a single RTX3090 GPU for 500 epochs with a batch size of 16. The inputs of CU-Reg are ultrasound slices with the size of 128$\times$128$\times$1$\times$1 and corresponding volumes with the size of 128$\times$128$\times$32$\times$1. In the 2D ultrasound frame branch, we employ a ResNet-34 [9] as the 2D encoder. In the training phase, the inputs are a current anchor frame and three adjacent frames, i.e., 128$\times$128$\times$1$\times$4. During inference, we manually adjust the input dimension by repeating the number of channels four times. The hyperparameter values are the optimal results obtained through ablation experiments.
@@ 3.1 Dataset and Evaluation Metrics
To evaluate the cardiac frame-to-volume registration network, a simulated dataset is generated by post-processing the public CAMUS dataset [13]. CAMUS contains 2D echocardiographic sequences with two- and four-chamber views of 500 patients, along with the masks of the left ventricular epicardium, these 2D sequences are expressed as 3D volumes in Cartesian coordinates with a unique grid resolution using the same interpolation procedure. For each original 3D volume, four transformations are generated by add random deviations to the identity transformation, the deviations of translation parameters of each transformation are within the range of 10 mm, and the deviations of the rotation parameters are within the range of 20 degrees. Based on these transformations, 2D slices and corresponding masks are sampled from the original volume. The sampled slices include 128$\times$128 pixels, and the pixel spacing is 0.62 mm×$\times$0.62 mm. Meanwhile, a new volume is sampled from the original volume based on the identity transformation with a volume size of 128$\times$128$\times$32 and a voxel spacing of 0.62 mm×$\times$0.62 mm×$\times$0.62 mm. In this way, we can obtain four volume-frame-mask pairs with true transformations for one original volume. All data are split at the patient level, with 3,600 volume-frame-mask pairs for training and 400 volume-frame-mask pairs for testing.
@@ 4.0.1 Acknowledgements
This work was supported in part by the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No.: T45-401/22-N), in part by the Hong Kong Innovation and Technology Fund under Grant GHP/080/20SZ, in part by the Innovation and Technology Fund under Guangdong-Hong Kong Technology Cooperation Funding Scheme under Grant GHP/050/20SZ, in part by the National Natural Science Foundation of China under Grant 62273328, and in part by the Guangdong-Hong Kong-Macao Research Team Project under Grant 2021B1515130003.
@@
The authors have no competing interests to declare.